The following equation describes a person’s optimum heart rate in beats per minute that should be achieved during exercise so that the exercise is most beneficial.
Heart rate = 0.6(220-a)
where a represents the age of the person in years.
Determine the optimum heart rate for a 35-year-old runner.
I figured out that the age of this runner is 22 years old, but can't figure out how to get the heartrate of a 35 year old.

Answer by stanbon(75887)

Heart rate = 0.6(220-a)
where a represents the age of the person in years.
Determine the optimum heart rate for a 35-year-old runner.
----------------------------
H(35) = 0.6(220-35)
= 0.6(185)
= 111 beats per minute
==============================
Cheers,
Stan H.